OpenStudy (anonymous):

2^5 x 2^-2=? First express the answer in exponential form. then evaluate the expression.

OpenStudy (gw2011):

(2^5)(2^-2)=(2^3)=(2)(2)(2)=8

OpenStudy (anonymous):

2^3 = 8?

OpenStudy (gw2011):

Another way of looking at the expression is: (2^5)(1/2)^2 since (2^-2) is equal to (1/2)^2 and you still get the same answer

OpenStudy (anonymous):

thanks.

OpenStudy (gw2011):

You're welcome